# Deep Multi-Task Learning Has Low Amortized Intrinsic Dimensionality

### v1

@@ Discussion.
Theorem 5.6 establishes that the multi-task generalization gap (the difference between empirical risk and expected risk) can be controlled by a term that expresses how compactly the models can be encoded. Because the inequality is uniform with respect to $E\in\mathcal{E}$, the multi-task encoding scheme can be chosen in a data-dependent way, where a tighter guarantee follows if the multi-task encoding scheme itself can be compactly represented.
To get a better intuition of this procedure, we consider two special cases. First, assume a naive encoding, where $\mathcal{E}$ consists of a single element, $E$, which itself is the naive multi-task model encoder that simply stores a $D$-dimensional parameter vector for each model in fixed-width form. Then $l(E)=0$ and $l_{E}(f_{1},\dots,f_{n})=O(nD)$, and the resulting complexity term is of the order $O(\sqrt{\frac{D}{m}})$, as classical VC theory suggests.
In contrast, assume the setting of Section 3, where each task is encoded as a weighted linear combination of $k$ dictionary elements, each of which is computed as the product of a large random matrix with an $l$-dimensional parameter vector. Setting $\mathcal{E}$ to be the set of all such bases, one obtains $l(E)=O(lk)$ (for now disregarding potential additional savings by storing coefficients more effectively than with fixed width bits), and $l_{E}(f_{1},\dots,f_{n})=O(nk)$. Consequently, the complexity term in Theorem 5.6 become $O(\sqrt{\frac{lk/n+k}{m}})$, i.e. the ambient dimension $D$ of the previous paragraph is replaced by exactly the amortized intrinsic dimension of Definition 3.1.
Table 2: Generalization guarantees (upper bound on test error, lower is better) for single-task and multi-task learning. For all tested datasets the multi-task analysis offers better guarantees than the single-task analysis, sometimes by a large margin. The bounds are also always non-vacuous. The fast-rate bound (Theorem 5.6) offers slightly stronger guarantees than the more elementary analysis of Theorem 5.6.
@@
Single task [%] 61.2 57.6 56.6 33.2 87.4 99.4
MTL (Theorem 5.6) [%] 23.0 40.4 39.4 22.2 52.9 86.9
MTL (Theorem 5.6) [%] 19.6 35.0 38.8 20.3 52.7 83.0
5.3 Proofs
In this section, we provide the proof for Theorem 5.6. The proof of Theorem 5.6 requires more involved tools, so we present it in the appendix.
Proof of Theorem 5.6.
@@ 5.3 Proofs
$\mathbb{P}\Bigg{\{}\exists E,f_{1},\dots,f_{n}:\;\mathcal{R}(f_{1},\dots,f_{n})-\widehat{\mathcal{R}}(f_{1},\dots,f_{n})$ (11)
$\quad\geq\sqrt{\frac{\big{(}l(E)+l_{E}(f_{1},\dots,f_{n})\big{)}\log(2)+\log\frac{1}{\delta}}{2mn}}\Bigg{\}}\leq\delta.$
By rearranging the terms, we obtain the claim of Theorem 5.6. ∎
5.4 Computing the Bounds Numerically
In order to apply Theorem 5.6 or Theorem 5.6 in the situation of Section 3, we use a meta-encoder that encodes the coefficient vectors, $v_{1},\dots,v_{k}$ of the basis representation (2), and a multi-task model encoder that jointly encodes the collection of per-task coefficients, $\alpha_{1},\dots,\alpha_{n}$ of the per-task representations 3.
Now, for any set of models $f_{1},\dots,f_{n}$ with global $v_{1},\dots,v_{k}$, and $\alpha_{1},\dots,\alpha_{n}$, the meta-encoder first stores the $r_{g}$ codebook entries as 16-bit floating-point values. Then, for each entry of the parameter vectors, it computes the index of the codebook entry that represents its value and it stores all indices together using arithmetic coding (Langdon, 1984), which exploits the global statistics of the index values. Analogously, the multi-task model encoder first stores the $l$ codebook entries, and then it represents the entries of $\alpha_{1},\dots,\alpha_{n}$ as indices in the code, and encodes them jointly using arithmetic coding. Note that while the latter process of encoding the model coefficients does not depend on the shared parameters, its decoding process does, because the actual network weights can only be recovered if also the representation basis, $Q$, is known.
The resulting guarantees then hold for the quantized networks. For Theorem 5.6, we can read off the guarantees on the multi-task risk directly from the explicit complexity term and the training error. For Theorem 5.6, we have to invert the $\operatorname{\mathbf{kl}}$-expression, which however is easily possible numerically. See Appendix A for more details.
5.4.1 Encoding all tasks together
[…]
5.5 Application to Transfer Learning
$\theta_{j}=\theta_{0}+Q\alpha+Pw,$ (12)
Theorem 5.3 yields generalization guarantees for this situation, in which only the encoding length of $\alpha$ and $w$ enters the complexity term, while $Q$ (like $P$) do not enter the bound. To encode the coefficients $\alpha$ and $w$, we can reuse the codebook of the MTL situation, which is now fixed and therefore does not have to be encoded or learn (and encode) a new one.
Dataset MNIST SP MNIST PL Folktables Products CIFAR10 CIFAR100
$m$ / $L$ 2000 / 10 2000 / 10 900/ 2 2000 / 2 453 / 3 450 / 10
No transfer [%] 60.4 58.1 54.8 26.1 87.1 98.2
Transfer from MTL [%] 15.1 49.7 40.1 16.1 58.0 84.8
6 Related Work
A seminal work in this area is Maurer (2006), where the author studies the feature learning formulation of MTL in which the systems learn a shared representation space and individual per-task classifiers within that representation. In the case of linear features and linear classifiers, he derives a generalization bounds based on Rademacher-complexity. Subsequently, a number of follow-up works extended and refined the results to cover other regularization schemes or complexity measures (Crammer & Mansour, 2012; Pontil & Maurer, 2013; Yousefi et al., 2018; Pentina & Lampert, 2017). However, none of these results are able to provide non-vacuous bounds for overparametrized models, such as deep networks.
A related branch of research targets the problems of representation or meta-learning. There, also multiple tasks are available for training, but the actual problem is to provide generalization guarantees for future learning tasks (Baxter, 2000). A number of works in this area rely on PAC-Bayesian generalization bounds (Pentina & Lampert, 2014; Amit & Meir, 2018; Rothfuss et al., ; Guan & Lu, 2022; Zakerinia et al., 2024), as we do for our Theorem 5.6. This makes them applicable to arbitrary model classes, including deep networks. These results, however, hold only under specific assumptions on the observed tasks, typically that these are themselves i.i.d. sample from a task environment. The complexity terms in their bounds are either not numerically computable, or vacuous by several orders of magnitude for deep networks, because they are computed in the ambient dimensions, not the intrinsic ones. Another difference to our work is that, even those parts of their bound that relate to multi-task generalization are sums of per-task contributions, which prevents the kind of synergetic effects that our joint encoding can provide.
7 Experimental Details
7.1 Datasets
MNIST Shuffled Pixels (SP): (Amit & Meir, 2018) each task is a random subset of the MNIST (LeCun, 1998) dataset in which 200 of the input pixels are randomly shuffled. The same shuffling is consistent across all samples of that task.
Folktables: (Ding et al., 2021) A tabular dataset consisting of public US census information. From personal features, represented in a binary encoding, the model should predict if a person’s income is above or below a threshold. Tasks correspond to different geographic regions.
7.2 Model Architectures
For the image datasets (MNIST-SP, MNIST-PL, split-CIFAR10, and split-CIFAR100) we use convolutional networks. For the MNIST experiments, we use the architecture from Amit & Meir (2018). For the CIFAR experiments, we use the network from Scott et al. (2024). For the vectorial dataset Product and the tabular dataset Folktables, we use 4-layer fully connected networks. For more details about the architectures and optimization procedure, see Appendix B.
The models’ ambient dimensions (number of network weights) range from approximately 12.000 to approximately 120.000, i.e. far more than the available number of samples per task. As random matrices $P$ for the single-task parametrization (1), we use the Kronecker product projector of Lotfi et al. (2022), $P=Q_{1}\otimes Q_{2}/\sqrt{D}$, for $Q_{1},Q_{2}\sim\mathcal{N}(0,1)^{\sqrt{D}\times\sqrt{d}}$. By this construction, the matrix $P\in\mathbb{R}^{D\times d}$ never has to be explicitly instantiated, which makes the memory and computational overhead tractable. For the multi-task representation (3), we use the analogous construction to form $Q^{\prime}=[P_{1},P_{2},\dots,P_{k}]=Q^{\prime}_{1}\otimes Q^{\prime}_{2}/\sqrt{D}\in\mathbb{R}^{D\times kl}$.
8 Conclusion
We presented a new parametrization for deep multi-task learning problems that directly allows one to read off its intrinsic dimension. We showed experimentally that for common benchmark tasks, the amortized intrinsic dimension per task can be much smaller than when training the same tasks separately. We derived two new encoding-based generalization bounds for multi-task learning, that result in non-vacuous guarantees on the multi-task risk for several standard benchmarks. To our knowledge, this makes them the first non-vacuous bounds for deep multi-task learning.
A limitation of our analysis in Section 3 is that it relies on a specific representation of the learned models. As such, the values we obtain for the amortized intrinsic dimensionality constitute only upper bounds to the actual, i.e. minimal, number of necessary degrees of freedom. In future work, we plan to explore the option of extending our results also to other popular multi-task parametrization, such as based on prototypes, MAML (Finn et al., 2017), or feature learning (Collobert & Weston, 2008). Note that our theoretical results in Section 5 readily apply to such settings, as long as we define suitable encoders. In this context, as well as in general, it would be interesting to study if our bounds could be improved further by exploiting more advanced techniques for post-training weight quantization (Rastegari et al., 2016; Frantar et al., 2023) or learning models directly in a quantized form (Hubara et al., 2018; Wang et al., 2023).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs
A.1 Comparison of the Bounds
Theorem 5.6 which in the PAC-Bayes literature is referred to as a fast-rate bound, can be tighter than the bound of the Theorem 5.6. Specifically, up to a different in the $\log$-term we can obtain the bound of the Theorem 5.6 by relaxing the bound of the Theorem 5.6. Based on Pinsker’s inequality, we have $2(p-q)^{2}\leq\operatorname{\mathbf{kl}}(q|p)$ and therefore
Combining Theorem 5.6 and equation (13) gives that for every $\delta>0$ with probability at least $1-\delta$, we have:
This is really similar to the bound of Theorem 5.6, except for the extra $\frac{\log 2\sqrt{mn}}{2mn}$, which is negligible for large $m$ or $n$.
Corollary A.1.
With the same setting as Theorem 5.6, with probability at least $1-\delta$, we have:
As Table 2 shows, the numeric bounds obtained this way can be tighter bound than the upper bound from the Theorem 5.6.
A.2 Proof of Theorem 5.6.
@@ A.1 Comparison of the Bounds
In this section, we provide the proof for Theorem 5.6. Since our hypothesis set is discrete, we use a union-bound approach, similar to the proof of Theorem 5.6. A similar result can be proved using the PAC-Bayes framework and change of measure, similar to the fast-rate bounds in Guan & Lu (2022), but we found our version to be conceptually simpler.
@@
Subsequently, we follow the steps of the proof of Theorem 5.6. Assume a fixed $\delta>0$. For any meta-encoder, $E$, and any tuple of models, $F=(f_{1},...,f_{n})$, we define a weight $w_{E;F}=\delta\cdot 2^{-l(E)}2^{-l_{E}(f_{1},\dots,f_{n})}$, where $l(\cdot)$ is the length function of the meta-encoder and $l_{E}(\cdot)$ is the length function of the encoder $E$. We instantiate (18) with a value $t_{E;F}$ such that $2\sqrt{mn}e^{-t_{E;F}mn}=w_{E;F}$, i.e. $t_{E;F}=-\frac{\log w_{E;F}}{mn}$. Therefore, we have with probability at least $1-w_{E,F}$:
@@
As mentioned earlier, Guan & Lu (2022) proved a fast-rate bound for meta-learning which consists of a multi-task bound and a meta bound. The multi-task bound provided here, shares a similar structure with the bound of Guan & Lu (2022), and a similar approach to upper-bound the MGF (Moment generating function). There are three main differences in how to use the bounds. The first one is that they approximate the upper-bound for the $\mathcal{R}(F)$ based on a closed-form approximation, which has poorer performance compared to the numerical optimization, and even in the cases where the empirical error is not small, their approximation can be even worse than the bound of the Theorem 5.6). The second difference is that they use Gaussian distributions on the networks that scale with ambient dimensionality, making the bounds vacuous by several orders of magnitude. The more structural difference is that the sum of the task-specific complexity terms appears in their bound. On the other hand, we have a joint complexity term for the task-specific part (which in our case is the length of the multi-task encoding), which as explained in the section 5.4.1, is much smaller than the sum of individual ones.
A.3 Auxiliary Lemmas
Lemma A.2.
@@ A.3 Auxiliary Lemmas
Lemma A.3.
(Maurer, 2004)[Theorem 1] Let $Y$ be the binomial random variable with distribution $Y\sim B\left(t,\frac{\mu}{t}\right)$. Then for the function $kl$ as defined in (7) we have:
@@
Lemma A.4.
@@
Let $f$ be the function $f(x)=mn\operatorname{\mathbf{kl}}(\frac{x}{mn}|\mathcal{R}(F))$, this function is convex. If we define $X_{i,j}=\ell_{i}({f_{i},z_{i,k}})$, since $f_{i}$s are fixed, and samples are independent, the random variables $X_{i,j}$s are independent. Therefore, $\sum X_{i,j}=mn\widehat{\mathcal{R}}(F)$, and $\operatorname*{\mathbb{E}}[\sum X_{i,j}]=mn\mathcal{R}(F)$. Let $Y\sim B(mn,\mathcal{R}(F))$. Because of Lemma A.2 we have
@@
Because of Lemma A.3, we have:

### v2

@@ Discussion.
Definition 4 establishes that the multi-task generalization gap (the difference between empirical risk and expected risk) can be controlled by a term that expresses how compactly the models can be encoded. Because the inequality is uniform with respect to $E\in\mathcal{E}$, the global parameter can be chosen in a data-dependent way, where tighter guarantees can be achieved if the global parameter itself can be compactly represented.
To get a better intuition of this procedure, we consider two special cases. First, assume a naive encoding, where no global shared parameter is encoded, and the multi-task model encoder simply stores a $D$-dimensional parameter vector for each model in fixed-width form. Then $l(E)=0$ and $l_{E}(f_{1},\dots,f_{n})=O(nD)$, and the resulting complexity term is of the order $O(\sqrt{\frac{D}{m}})$, as classical VC theory suggests.
In contrast, assume the setting of Section 3, where each task is encoded as a weighted linear combination of $k$ dictionary elements, each of which is computed as the product of a large random matrix with an $l$-dimensional parameter vector. Setting $\mathcal{E}$ to be the set of all such bases, one obtains $l(E)=O(lk)$ (for now disregarding potential additional savings by storing coefficients more effectively than with fixed width bits), and $l_{E}(f_{1},\dots,f_{n})=O(nk)$. Consequently, the complexity term in Theorem 4 become $O(\sqrt{\frac{lk/n+k}{m}})$, i.e. the ambient dimension $D$ of the previous paragraph is replaced by exactly the amortized intrinsic dimension of Definition 1.
Table 2: Generalization guarantees (upper bound on test error, lower is better) for single-task and multi-task learning. For all tested datasets the multi-task analysis offers better guarantees than the single-task analysis, sometimes by a large margin. The bounds are also always non-vacuous. The fast-rate bound (Definition 4) offers improved guarantees compared to the more elementary Definition 4.
@@
model ConvNet ConvNet MLP MLP ConvNet ViT ConvNet ViT
Single task 0.612 0.576 0.566 0.332 0.874 0.660 0.994 0.905
MTL (Definition 4) 0.230 0.404 0.394 0.222 0.529 0.319 0.869 0.665
MTL (Definition 4) 0.196 0.350 0.388 0.203 0.527 0.280 0.830 0.658
5.3 Computing the bounds numerically
In order to apply Definition 4 or Definition 4 in the situation of Section 3, we use a meta-encoder that encodes the coefficient vectors, $v_{1},\dots,v_{k}$ of the basis representation (2), and a multi-task model encoder that jointly encodes the collection of per-task coefficients, $\alpha_{1},\dots,\alpha_{n}$ of the per-task representations 3.
Now, for any set of models $f_{1},\dots,f_{n}$ with global $v_{1},\dots,v_{k}$, and $\alpha_{1},\dots,\alpha_{n}$, the meta-encoder first stores the $r_{g}$ codebook entries as 16-bit floating-point values. Then, for each entry of the parameter vectors, it computes the index of the codebook entry that represents its value and it stores all indices together using arithmetic coding (Langdon, 1984), which exploits the global statistics of the index values. Analogously, the multi-task model encoder first stores the $l$ codebook entries. It then represents the entries of $\alpha_{1},\dots,\alpha_{n}$ as indices in the code, and encodes them jointly using arithmetic coding. Note that while the latter process of encoding the model coefficients does not depend on the shared parameters, its decoding process does, because the actual network weights can only be recovered if also the representation basis, $Q$, is known.
The resulting guarantees then hold for the quantized networks. For Definition 4, we can read off the guarantees on the multi-task risk directly from the explicit complexity term and the training error. For Definition 4, we have to invert the $\operatorname{\mathbf{kl}}$-expression, which is easily possible numerically. See Appendix A for more details.
5.3.1 Encoding all tasks together
[…]
5.4 Application to transfer learning
model ConvNet ConvNet MLP MLP ConvNet ViT ConvNet ViT
No transfer 0.604 0.581 0.548 0.261 0.871 0.700 0.982 0.911
MTL + Transfer 0.151 0.497 0.401 0.161 0.580 0.283 0.848 0.606
$\theta_{j}=\theta_{0}+Q\alpha+Pw,$ (7)
Theorem 1 yields generalization guarantees for this situation, in which only the encoding length of $\alpha$ and $w$ enters the complexity term, while $Q$ (like $P$) do not enter the bound. To encode the coefficients $\alpha$ and $w$, we can reuse the codebook of the MTL situation, which is now fixed and therefore does not have to be encoded or learn (and encode) a new one.
6 Related work
A seminal work in this area is Maurer (2006), where the author studies the feature learning formulation of MTL in which the systems learn a shared representation space and individual per-task classifiers within that representation. In the case of linear features and linear classifiers, he derives a generalization bounds based on Rademacher-complexity. Subsequently, a number of follow-up works extended and refined the results to cover other regularization schemes or complexity measures (Crammer & Mansour, 2012; Pontil & Maurer, 2013; Yousefi et al., 2018; Pentina & Lampert, 2017; Du et al., 2021). However, none of these results are able to provide non-vacuous bounds for overparametrized models, such as deep networks.
A related branch of research targets the problems of representation or meta-learning. There, also multiple tasks are available for training, but the actual problem is to provide generalization guarantees for future learning tasks (Baxter, 2000) A number of works in this area rely on PAC-Bayesian generalization bounds (Pentina & Lampert, 2014; Amit & Meir, 2018; Liu et al., 2021; Guan & Lu, 2022; Riou et al., 2023; Friedman & Meir, 2023; Rezazadeh, 2022; Rothfuss et al., 2023; Ding et al., 2021b; Tian & Yu, 2023; Farid & Majumdar, 2021; Zakerinia et al., 2024), as we do for our Definition 4. This makes them applicable to arbitrary model classes, including deep networks. These results, however, hold only under specific assumptions on the observed tasks, typically that these are themselves i.i.d. sample from a task environment. The complexity terms in their bounds are either not numerically computable, or vacuous by several orders of magnitude for deep networks, because they are computed in the ambient dimensions, not the intrinsic ones. Another difference to our work is that, even those parts of their bound that relate to multi-task generalization are sums of per-task contributions, which prevents the kind of synergetic effects that our joint encoding can provide.
7 Conclusion
We presented a new parametrization for deep multi-task learning problems that directly allows one to read off its intrinsic dimension. We showed experimentally that for common benchmark tasks, the amortized intrinsic dimension per task can be much smaller than when training the same tasks separately. We derived two new encoding-based generalization bounds for multi-task learning, that result in non-vacuous guarantees on the multi-task risk for several standard datasets and model architectures. To our knowledge, this makes them the first non-vacuous bounds for deep multi-task learning.
A limitation of our analysis in Section 3 is that it relies on a specific representation of the learned models. As such, the values we obtain for the amortized intrinsic dimensionality constitute only upper bounds to the actual, i.e. minimal, number of necessary degrees of freedom. In future work, we plan to explore the option of extending our results also to other popular multi-task parametrization, such as methods based on prototypes, MAML (Finn et al., 2017), or feature learning (Collobert & Weston, 2008). Note that our theoretical results in Section 5 readily apply to such settings, as long as we define suitable encoders. In this context, as well as in general, it would be interesting to study if our bounds could be improved further by exploiting more advanced techniques for post-training weight quantization (Rastegari et al., 2016; Frantar et al., 2023) or learning models directly in a quantized form (Hubara et al., 2018; Wang et al., 2023).
Acknowledgements
This research was supported by the Scientific Service Units (SSU) of ISTA through resources provided by Scientific Computing (SciComp).
Appendix A Proofs
A.1 Comparison of the bounds
Definition 4 which in the PAC-Bayes literature is referred to as a fast-rate bound, can be tighter than the bound of the Definition 4. Specifically, up to a different in the $\log$-term we can obtain the bound of the Definition 4 by relaxing the bound of the Definition 4. Based on Pinsker’s inequality, we have $2(p-q)^{2}\leq\operatorname{\mathbf{kl}}(q|p)$ and therefore
Combining Definition 4 and equation (8) gives that for every $\delta>0$ with probability at least $1-\delta$, we have:
This is very similar to the bound of Definition 4, differing by the additional term $\frac{\log 2\sqrt{mn}}{2mn}$, which is negligible for large $m$ or $n$.
Corollary 2.
With the same setting as Definition 4, with probability at least $1-\delta$, we have:
As Table 2 shows, the numeric bounds obtained this way can be tighter bound than the upper bound from the Definition 4.
A.2 Proof of Theorem 4
[…]
Proof.
@@ A.2 Proof of Theorem 4
$\mathbb{P}\Bigg{\{}\exists E,F:\;\mathcal{R}(F)-\widehat{\mathcal{R}}(F)$ $\geq\sqrt{\frac{\Big{(}l(E)+l_{E}(F)\Big{)}\log(2)+\log\frac{1}{\delta}}{2mn}}\Bigg{\}}\leq\delta.$ (15)
By rearranging the terms, we obtain the claim of Definition 4. ∎
A.3 Proof of Theorem 4.
@@
In this section, we provide the proof for Definition 4. Since our hypothesis set is discrete, we use a union-bound approach, similar to the proof of Definition 4. A similar result can be proved using the PAC-Bayes framework and change of measure, similar to the fast-rate bounds in Guan & Lu (2022), but we found our version to be conceptually simpler.
@@
Subsequently, we follow the steps of the proof of Definition 4. Assume a fixed $\delta>0$. For any meta-encoder, $E$, and any tuple of models, $F=(f_{1},...,f_{n})$, we define a weight $w_{E;F}=\delta\cdot 2^{-l(E)}2^{-l_{E}(f_{1},\dots,f_{n})}$, where $l(\cdot)$ is the length function of the meta-encoder and $l_{E}(\cdot)$ is the length function of the encoder $E$. We instantiate (17) with a value $t_{E;F}$ such that $2\sqrt{mn}e^{-t_{E;F}mn}=w_{E;F}$, i.e. $t_{E;F}=-\frac{\log w_{E;F}}{mn}$. Therefore, we have with probability at least $1-w_{E,F}$:
@@
As mentioned earlier, Guan & Lu (2022) proved a fast-rate bound for meta-learning which consists of a multi-task bound and a meta bound. The multi-task bound provided here, shares a similar structure with the bound of Guan & Lu (2022), and a similar approach to upper-bound the MGF (Moment generating function). There are three main differences in how to use the bounds. The first one is that they approximate the upper-bound for the $\mathcal{R}(F)$ based on a closed-form approximation, which has poorer performance compared to the numerical optimization, and even in the cases where the empirical error is not small, their approximation can be even worse than the bound of the Definition 4). The second difference is that they use Gaussian distributions on the networks that scale with ambient dimensionality, making the bounds vacuous by several orders of magnitude. The more structural difference is that the sum of the task-specific complexity terms appears in their bound. On the other hand, we have a joint complexity term for the task-specific part (which in our case is the length of the multi-task encoding), which as explained in the section 5.3.1, is much smaller than the sum of individual ones.
A.4 Auxiliary lemmas
Lemma 3.
@@ A.4 Auxiliary lemmas
Lemma 4.
(Maurer, 2004)[Theorem 1] Let $Y$ be the binomial random variable with distribution $Y\sim B\left(t,\frac{\mu}{t}\right)$. Then we have:
@@
where $\operatorname{\mathbf{kl}}(q|p)=q\log\frac{q}{p}+(1-q)\log\frac{1-q}{1-p}$ is the Kullback-Leibler divergence between Bernoulli distributions with mean $q$ and $p$.
Lemma 5.
@@
Let $f$ be the function $f(x)=mn\operatorname{\mathbf{kl}}(\frac{x}{mn}|\mathcal{R}(F))$, this function is convex. If we define $X_{i,j}=\ell_{i}({f_{i},z_{i,k}})$, since $f_{i}$s are fixed, and samples are independent, the random variables $X_{i,j}$s are independent. Therefore, $\sum X_{i,j}=mn\widehat{\mathcal{R}}(F)$, and $\operatorname*{\mathbb{E}}[\sum X_{i,j}]=mn\mathcal{R}(F)$. Let $Y\sim B(mn,\mathcal{R}(F))$. Because of Lemma 3 we have
@@
Because of Lemma 4, we have: